HUBBLE SPACE TELESCOPE DAILY REPORT #2169 PERIOD COVERED: 0000Z (UTC) 07/20/98 - 0000Z (UTC) 07/21/98 Daily Status Report as of 202/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOC/96/WF/PC-2 7274 (Secular Changes in the Jet of M-87) The Faint Object Camera (f/96) and the WF/PC-2 were both used to make observations of the secular changes in the jet coming off of the active galactic nuclei of galaxy M-87. Past observations with the FOC have shown changes as small as 0.03", and provide a unique opportunity to study proper motions and structural changes of the jet. An internal flat calibration image was also taken. The observations were completed as scheduled, and no anomalies were reported. 1.2 Completed NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. The observations were completed as planned, and no anomalies were reported. 1.3 Completed Three Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed NIC/2 7887 (A Survey of Gravitational Lenses as Cosmological Tools II) The Near Infrared Camera (Camera Two) was used to make observations of the gravitational lens candidate system SBS1520+53. The propose is to extend our IR survey of gravitational lenses and gravitational lens candidates by obtaining NICMOS images of 14 additional lens systems or candidates. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed Two Sets of STIS/CCD 8064 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal - Backup) The Space Telescope Imaging Spectrograph (CCD) was used to conduct a non-scripted parallel POMS test. The observations were completed as scheduled, and no problems were reported. 1.6 Completed Two Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.7 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed Two Sets of WF/PC-2 8061 (POMS Test Proposal: WF-II Backup Parallel Archive Proposal) This is a POMS test proposal designed to simulate scientific plans. The proposal was executed as planned, and no anomalies were reported. 1.9 Completed NIC/1/2 7329 (The Nature of the Damped LyAlpha Absorbers -- A New Study of Young Galaxies) The Near Infrared Camera (Cameras One and Two) was used to make observations of the young galaxy Q1759+7539 in order to study the nature of its damped Lyman-Alpha Absorbers (DLAS). The damped Lyman-Alpha absorption line systems, occasionally found in quasar spectra, provide one of only a few practical methods of detecting and studying significant numbers of galaxies at high redshifts. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed NIC/2/STIS/CCD/WF/PC-2 7220 (Imaging of Quasar Host Galaxies) The Near Infrared Camera (Camera Two), the Space Telescope Imaging Spectrograph (CCD), and the WF/PC-2 were used to make direct imaging observations of the host galaxy for the broad absorption line quasar PG1700+518. The observations were completed as planned, and no anomalies were reported. 1.11 Completed WF/PC-2 7386 (A Snapshot Survey of Nova Shells) The WF/PC-2 was used to make observational images of the super nova shell V1974-Cyg in order to detect and measure its expanding gaseous shell. These observations will trace the distribution of hot and cool gas in the nova shells, and show their degree of inhomogeneity on both large and small scales. The observations were completed as scheduled, and no problems were noted. 1.12 Completed NIC/1/2 7285 (Collimation of Astrophysical Jets: The Proto-Planetary Nebula He 3-1475) The Faint Object Camera (f/96) was used to make observations of proto-planetary nebula HE-1475 in order to get a collimation of its astrophysical jets. Recent observations of this proto-planetary nebula in the N II Lambda 6584 A0 line with the HST WF/PC-2 have revealed what appear the be large scale flows being collimated into narrow bipolar jets. The observations were completed as scheduled, and no anomalies were reported. 1.13 Completed NIC/1/2/3 7824 (Deep NIC Images of 20 High-Z Damped Ly-Alpha Galaxies) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the high-Z damped Lyman-Alpha (DLA) galaxy PKS0528-25. The HST imaging together with ground-based spectroscopy will unify absorption-line studies and faint galaxy imaging studies, providing a comprehensive picture of the history of assembly, star formation, gas consumption, and chemical enrichment in the DLA galaxies. The observations were completed as planned, and no problems were reported. 1.14 Completed STIS/CCD 7910 (STIS Extragalactic Parallel Survey) The Space Telescope Imaging Spectrograph (CCD) was used to make extragalactic slitless spectrographic observations as part of the parallel high galactic latitude survey to search for emission line galaxies. Data was taken in parallel, and are not subject to proprietary time. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 9 Successful : 9 Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled : 16 Successful : 16 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC eror counter was cleared two times. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH